(a) The Cartesian coordinates of a point in the xy-plane are (x, y) = (-3.31, -2.49) m. Find the polar coordinates of this point.
r =
m
θ =
°

(b) Convert (r, θ) = (4.87 m, 35.5°) to rectangular coordinates.
x =
m
y =
m
1
Expert's answer
2019-09-05T09:09:49-0400

a)


"r=\sqrt{(-3.31)^2+(-2.49)^2}=4.14\ m"

"\theta=180+\arctan{\frac{2.49}{3.31}}=180+37=217\degree"

b)


"x=4.87\cos{35.5}=3.97\ m"

"y=4.87\sin{35.5}=2.83\ m"